an accererlating force of 500N changes an object from 40 m/s to 60 m/s in a distance of 100m. the mass is 25 kg determine the force of friction acting opposite this force.
1
Expert's answer
2011-10-25T08:06:23-0400
Let's mark the fricting force by F. Using the second Newton's law we obtain:
500-F = 25a (1)
The average speed of an object is
(60+40)/2 = 50 m/s.
So, moving time T:
T = 100/50 = 2 s.
Let's find the acceleration of an object:
a = (60-40)/T = 20/2 = 10 m/s^2.
From (1) we obtain:
F = 500-25a = 500-25*10 = 250 N.
So, the friction force has value 250N.